An iron block, of mass 10 Kg, rests on a wooden plane inclined at 300 to the horizontal. It is found that the least force parallel to the plane which causes the block to slide up the plane is 10 N. Calculate the coefficient of sliding friction between wood and iron.
e
1
Expert's answer
2020-11-05T04:07:55-0500

here given angle = 300 of inclination, on incline plane if angle given is "" \alpha "" and coefficient of friction is "k" then, sliding friction force = "k\times m \times g \times Sin(\alpha)"

here in our question this Force will be equal to 10N, Mass given 10Kg, g=gravity =9.8 m/s2...

Hence ""k \times m\times g\times Sin(\alpha) = 10""

=> "k= coefficient of friction = \frac{10}{m \times g \times sin(\alpha)}"

=> "k = \frac{10}{10\times 9.8\times 0.5} = \frac{1}{4.9} = 0.204"

Hence Answer = 0.2